Prove the following result:

A function f that is decreesing on [a,b] is integrable on [a,b].


1
Expert's answer
2021-03-12T11:04:22-0500

Given that f is decreasing on [a,b]. Thus, "f(b) \leq f(x) \leq f(a)" and f is bounded on [a,b]. Given "\epsilon >0 \, \, \exist k >0 \ni k[f(a)-f(b)]< \epsilon" .


Let P = "\{x_0, x_1,..., x_n\} \ni ∆x_i \leq k" be a partition on [a,b].

Since f is decreasing it follows that "m_i = f(x_i)" and "M_i = f(x_{i-1}) i =1,2,...,n"

where "m_i" is the greatest lower bound of f on "[x_{i-1},x_i]" and "M_i" is the lowest upper bound of f on the interval "[x_{i-1},x_i]" "U(f,P) - L(f,P) = \sum_{i=1}^{n}[f(x_{i-1})-f(x_i)]∆x_i \leq k\sum_{i=1}^{n}[f(x_{i-1})-f(x_i)]" = "k[f(x_0)-f(x_n)]=k[f(a)-f(b)]< \epsilon"


Hence f is integrable on [a,b].